OpenStudy (anonymous):

k/6+2=0 solve each equation

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

k/6+2=0 k/6+2-2=-2 k/6=-2 6*k/6=-2*6 k=-12